# Computing almost commuting bases of ODOs and Gelfand-Dikey hierarchies

### v1

@@ 3.1 Formal differential operators
Then $\mathcal{R}$ is a differential ring whose ring of constants equals the $\mathbf{C}$. We call $\mathcal{R}[\partial]$ the ring of formal differential operators in the differential variables $U$ and its elements formal differential operators.
Weight function. We define a weight function $w$ over $\mathcal{R}$ given by $w(u_{\ell})=\ell$, and
$w(u_{\ell}^{(k)})=\ell+q,\ k=1,2,\dots.$ (6)
As in [18, Proposition 4.2], we call an operator $P\in\mathcal{R}[\partial]$ homogeneous of weight $r$ with respect to $w$ if the coefficient of $\partial^{i}$ in $P$ has weight $r-i$. This is equivalent to extending $w$ to $\mathcal{R}[\partial]$ by setting $w(\partial)=1$. From the homogeneity of the commutation rule $\partial p=p\partial+\partial(p)$, for every $p\in\mathcal{R}$ it follows that the product of two operators that are homogeneous of weights $r$ and $s$ is homogeneous of weight $r+s$.
@@
which is homogeneous of weight $n$. The next theorem is our revision of [21, Proposition 2.4], which establishes the foundation of the contributions of this paper.
@@
We call a pseudo-differential operator in $\mathcal{R}((\partial^{-1}))$ homogeneous of weight $r$ if the coefficient of $\partial^{i}$ is homogeneous of weight $r-i$, with $i\in\mathbb{Z}$. It also holds in $\mathcal{R}((\partial^{-1}))$ that the product of two operators that are homogeneous of weights $r$ and $s$ is homogeneous of weight $r+s$. It follows immediately that if a pseudo-differential operator is homogeneous of weight $r$ then its positive part is also homogeneous of weight $r$.
By [18, Proposition 4.2], the unique monic $n$-th root $Q=L^{1/n}$ of $L$ is a pseudo-differential operator in $\mathcal{R}((\partial^{-1}))$ and it is homogeneous of weight $1$. Therefore $Q^{j}$ is homogeneous of weight $j$ and we conclude that $P_{j}=(Q^{j})_{+}$ is homogeneous of weight $j$. Since $L$ is monic and in normal form then so is $Q$, which implies that each $P_{j}$ is monic and in normal form.
To prove its uniqueness, let us consider any other $\mathbf{C}$-basis $\mathcal{T}=\{T_{j}\mid j\in\mathbb{N}\}$ of $W(L)$ satisfying the homogeneity condition. It is necessary that $T_{j}=c_{j}P_{j}$, with $c_{j}\in\mathbf{C}$, since otherwise it would not be homogeneous anymore. Since we also require the basis to be monic, we can conclude that $c_{j}=1$ for all $j\in\mathbb{N}$, proving $\mathcal{T}=\mathcal{H}(L)$. ∎
@@
is a $\mathbb{Q}$-basis of $W_{M}(L)$. Our goal is to compute the basis $\mathcal{H}_{M}(L)$.
@@
Let us illustrate the homogeneous almost commuting basis for the case $n=3$, that leads to the Boussinesq hierarchy, which will be studied in Section 5.1. Consider a formal third-order differential operator
@@
Its cubic root is a pseudo differential operator in $\mathcal{R}((\partial^{-1}))$
@@
[…]
In Section 4 we provide a deterministic algorithm to compute the sequence of $P_{j}$ defined in (8), up to a fixed bound $M$ for any generic differential operator of order $n$ as in (7). Our philosophy is to avoid computations in the ring of pseudo-differential operators $\mathcal{R}((\partial^{-1}))$, since the formal implementation of $\mathcal{R}((\partial^{-1}))$ in a computational algebra system is not available so far.
@@ 4 Basis of almost-commuting ODOs
Extended weight function. Let us define a new set of differential variables $Y=\{y_{2},\dots,y_{M}\}$ over $\mathcal{R}$. Let $Y_{m}$ denote $\{y_{2},\ldots,y_{m}\}$ for $m=2,\ldots,M$. Extending the derivation in $\mathcal{R}$ formally by
@@
which is homogeneous of weight $m$. Consequently, the Lie bracket of $L$ and $\tilde{P}_{m}$ is homogeneous of weight $n+m$ in $\mathbb{Q}\{Y_{m}\}[\partial]$. By Lemma 2.1, the Lie bracket $[L,\tilde{P}_{m}]$ provides a formal differential operator of order $n+m-3$ since both operators are in normal form. The next equation will be of importance in the remaining parts of this paper:
$[L,\tilde{P}_{m}]=E_{m,0}(U,Y_{m})+E_{m,1}(U,Y_{m})\partial+\ldots+E_{m,n+m-3}(U,Y_{m})\partial^{n+m-3},$ (15)
@@
$\epsilon_{Z}:\mathcal{R}\{Y_{m}\}\rightarrow\mathcal{R}\ ,\ {\varepsilon_{Z}(y_{\ell}^{(k)})=q_{\ell}^{(k)}},$ (17)
and denote $f(U,Z):=\epsilon_{Z}(f(U,Y_{m}))$ Abusing notation, we define the extension of the previous morphism to the corresponding rings of differential operators
@@
Let us fix a positive integer $m\geq 2$, and the homogeneous operator $P_{m}$ in the homogeneous basis $\mathcal{H}(L)$ of $W(L)$ (see Theorem 3.6). Let us consider $Z=\{q_{2},\dots,q_{m}\}\subset\mathcal{R}$ with weights $w(q_{\ell})=\ell$. The following statements are equivalent:
@@
    $Z$ is a solution of system $\mathbb{E}_{m,n}$, that is
@@
The result will follow from the next observation
$[L,\epsilon_{Z}(\tilde{P}_{m})]=\epsilon_{Z}([L,\tilde{P}_{m}])=\sum_{j=3}^{n+m}E_{m,n+m-j}(U,Z)\partial^{n+m-j}.$ (19)
If we assume 1, since $[L,P_{m}]$ is an operator of order $n-2$, then $Z$ is a solution of system (16). Now assume 2. Then, by (19), $\epsilon_{Z}(\tilde{P}_{m})$ almost commutes with $L$ and it is an homogeneous operator of weight $m$, by the weight condition on $q_{\ell}$. Therefore, by the uniqueness in Theorem 3.6, the operator $\epsilon_{Z}(\tilde{P}_{m})$ equals $P_{m}$. ∎
@@
For each $n,m\in\mathbb{N}$, $n,m\geq 2$, the system $\mathbb{E}_{m,n}$ has a unique solution $Z=\{q_{2},\dots,q_{m}\}\in\mathbb{Q}\{U\}$ in the variables $Y$, with weights $w(q_{\ell})=\ell$. We call $Z$ the weighted solution of $\mathbb{E}_{m,n}$
@@ 2 LB $\leftarrow[L,\tilde{P}_{m}]$;
3 $\mathbb{E}_{m,n}$$\leftarrow\left\{\textrm{coeff }(\texttt{LB},\partial^{i})\ :\ i\in\{n-1,\ldots,n+m-3\}\right\}$;
@@ 4 Compute the unique weighted solution $Z$ of system $\mathbb{E}_{m,n}$ in (16);
The previous results induce Algorithm 1 to compute the almost commuting operator $P_{m}$ of the basis $\mathcal{H}(L)$. Iterating Algorithm 1 for $m=2,\ldots,M$ and setting $P_{1}=\partial$, we can compute the homogeneous basis of $W_{M}(L)$. We will discuss step 4 of this algorithm in the remaining parts of this section.
@@ 4.1 Triangular differential system
is a triangular system in the set of differential variables $Y=\{y_{2},\ldots,y_{m}\}$. In addition, each differential polynomial $E_{m,m+n-i}(U,Y_{m})$ will be proved to be homogeneous of weight $i$.
@@
$\tilde{P}_{\ell+1}=\tilde{P}_{\ell}\partial+y_{\ell+1},\,\,\,\ell\geq 1,\,\,\textrm{ defining }\tilde{P}_{1}:=\partial.$ (20)
This induces the following recursive formula to compute the Lie-brackets $[L,\tilde{P}_{m}]$.
@@
Let $L$ as in (7). Then, for $\ell\geq 1$
$[L,\tilde{P}_{\ell+1}]=[L,\tilde{P}_{\ell}]\partial+\tilde{P}_{\ell}[L,\partial]+[L,y_{\ell+1}].$ (21)
@@
We can see that the first summand adds the coefficients w.r.t. $\partial^{i}$ of $[L,\tilde{P}_{\ell}]$ as coefficients of $\partial^{i+1}$ in $[L,\tilde{P}_{\ell+1}]$. The second summand can be easily computed as a product of differential operators (see Lemma 4.3). Finally, the third summand has also a very generic formula (see Lemma 4.4) and is the only summand that involves the differential variable $y_{\ell+1}$.
@@
    $[L,\partial]=-\left(u_{n}^{\prime}+u_{n-1}^{\prime}\partial+\ldots+u_{2}^{\prime}\partial^{n-2}\right)$. Therefore, it is a homogeneous operator of weight $n+1$.
@@
    For $\ell\geq 2$, the differential operator $Q_{\ell}:=\tilde{P}_{\ell}[L,\partial]$ has order $n+\ell-2$, and it is a homogeneous operator of weight $\ell+n+1$. In particular, the coefficients of $Q_{\ell}$ only contain the differential variables $y_{2},\ldots,y_{\ell}$ with no derivatives. Moreover, the coefficient of $y_{\ell-i}\partial^{k}$ in $Q_{\ell}$ is 00, for $k>n+i-2$.
@@
The formula in 1 follows from the commuting rule $\partial u_{k}=u_{k}\partial+u_{k}^{\prime}$. Next, for proving 1, we first observe that $\rm ord(Q_{\ell})=\ell+n-2$, and, since $\tilde{P}_{\ell}$ is homogeneous of weight $\ell$ and $[L,\partial]$ is homogeneous of weight $n+1$, then $Q_{\ell}$ is also homogeneous of weight $\ell+n+1$. Moreover, we can compute
@@
$=-u_{2}^{\prime}\partial^{n+\ell-2}-\left(\ell u_{2}^{\prime\prime}+u_{3}^{\prime}\right)\partial^{n+\ell-3}-\sum_{i=0}^{\ell-2}y_{\ell-i}H_{n+i-2}+\left(\textrm{l.o.t in }\mathbf{C}\{U\}[\partial]\right),$
with $H_{n+i-2}$ a ordinary differential operator in $\mathbf{C}\{U\}[\partial]$ of order at most $n+i-2$. Then, the result follows. ∎
@@
Let $\ell$ be a positive integer. Then $[L,y_{\ell+1}]$ is a homogeneous differential operator of order $n-1$, weight $n+\ell+1$, and leading coefficient $ny_{\ell+1}^{\prime}$.
@@
Computing the commutator $[L,y_{\ell+1}]=Ly_{\ell+1}-y_{\ell+1}L$, we obtain
$\hss\begin{array}[]{rl}[L,y_{\ell+1}]=&\displaystyle\left(\partial^{n}+\sum_{i=0}^{n-2}u_{n-i}\partial^{i}\right)y_{\ell+1}-\left(y_{\ell+1}\partial^{n}+\sum_{i=0}^{n-2}y_{\ell+1}u_{n-i}\partial^{i}\right)\\ =&\displaystyle\left(\sum_{k=0}^{n}\binom{n}{k}y_{\ell+1}^{(n-k)}\partial^{k}+\sum_{i=0}^{n-2}\sum_{k=0}^{i}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\partial^{k}\right)-\left(y_{\ell+1}\partial^{n}+\sum_{k=0}^{n-2}y_{\ell+1}u_{n-i}\right)\\ =&ny_{\ell+1}^{\prime}\partial^{n-1}+\displaystyle\sum_{k=0}^{n-2}\binom{n}{k}y_{\ell+1}^{(n-k)}\partial^{k}+\sum_{i=0}^{n-2}\sum_{k=0}^{i}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\partial^{k}-\sum_{k=0}^{n-2}y_{\ell+1}u_{n-i}\partial^{k}\\ =&ny_{\ell+1}^{\prime}\partial^{n-1}+\displaystyle\sum_{k=0}^{n-2}\binom{n}{k}y_{\ell+1}^{(n-k)}\partial^{k}+\sum_{k=2}^{n-2}\sum_{i=k}^{n-2}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\partial^{k}-\sum_{k=0}^{n-2}y_{\ell+1}u_{n-i}\partial^{k}\\ =&ny_{\ell+1}^{\prime}\partial^{n-1}+\displaystyle\sum_{k=0}^{n-2}\left(\binom{n}{k}y_{\ell+1}^{(n-k)}+\sum_{i=k}^{n-2}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}-y_{\ell+1}u_{n-i}\right)\partial^{k}\\ =&ny_{\ell+1}^{\prime}\partial^{n-1}+\displaystyle\sum_{k=0}^{n-2}\left(\binom{n}{k}y_{\ell+1}^{(n-k)}+\sum_{i=k+1}^{n-2}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\right)\partial^{k}\\ =&ny_{\ell+1}^{\prime}\partial^{n-1}+\textrm{l.o.t.}\end{array}\hss$
@@
We establish next the triangularity of the system $\mathbb{E}_{m,n}$ defined in (16), and we provide algorithms to solve it afterwards.
@@
with $e_{m,i-2}\in\mathcal{R}\{Y_{i-2}\}$ homogeneous of weight $i$ for $i\geq 3$ (recall that $Y_{1}=\emptyset$). More precisely, $E_{m,n+m-3}=ny^{\prime}_{2}-mu^{\prime}_{2}$.
@@
We proceed by induction on $m$. First, for $m=2$ we use Lemmas 4.2 and 4.3 to we obtain
$[L,\tilde{P}_{2}]=-2\sum_{p=1}^{n-1}u^{\prime}_{p+1}\partial^{n-p}-\sum_{p=2}^{n}u^{\prime\prime}_{p}\partial^{n-p}+[L,y_{2}].$
@@
Now, assume the theorem holds for $m\geq 2$ and let us prove the case $m+1$. Using the recursive structure of $[L,\tilde{P}_{m+1}]$ from Lemma 4.2, formula (14) and Lemma 4.3, the following recursion holds
$[L,\tilde{P}_{m+1}]=\sum_{j=0}^{n+m-3}E_{m,j}\partial^{j+1}-\textrm{\small{$\left(\partial^{m}+\sum_{j=0}^{m-2}y_{m-j}\partial^{j}\right)\left(\sum_{p=2}^{n}u^{\prime}_{p}\partial^{n-p}\right)$}}+[L,y_{m+1}].$
We use the following notations to simplify the previous equality:
$\left\{\begin{array}[]{rl}A_{j}:=&\partial^{j}\displaystyle\sum_{p=2}^{n}u^{\prime}_{p}\partial^{n-p}\in\mathcal{R}[\partial]\ ,\textrm{ of order }n+j-2,\\ \alpha_{j,\ell}:=&\textrm{coeff }(A_{j},\partial^{\ell})\in\mathcal{R},\vspace{1em}\\ R_{m+1}:=&[L,y_{m+1}]-ny^{\prime}_{m+1}\partial^{n-1}\in\mathcal{R}\{y_{m+1}\}[\partial]\textrm{ of order }n-2.\end{array}\right.$
@@
$[L,\tilde{P}_{m+1}]=\sum_{j=0}^{n+m-3}E_{m,j}\partial^{j+1}-A_{m}-\sum_{j=0}^{m-2}y_{m-j}A_{j}+ny^{\prime}_{m+1}\partial^{n-1}+R_{m+1}.$
@@
Since $A_{j}$ has order $n+j-2$, it is clear that $\alpha_{j,\ell}=0$ for $\ell>n+j-2$, so, in particular, $\alpha_{j,n+m-2}=0$ for $j<m$. Moreover, we can observe that $\alpha_{j,n+j-2}=u_{2}^{\prime}$. Hence, using the recursion for $E_{m,n+m-3}$ we get:
@@
Finally, for all the intermediate equations, $14\leq\ell\leq m+$ 1, we observe that
@@
But, by induction hypothesis, the first summand equals $E_{m,n+m-\ell}=ny_{\ell-1}^{\prime}+e_{m,\ell-2}$, where $e_{m,\ell-2}$ is in $\mathcal{R}\{Y_{i-2}\}$. The second summand is in $\mathcal{R}$, and we can observe that the third summand involve the variables $y_{k}$ for $k=2,\ldots,\ell-3$. Observe that in the case of $\ell=4$ this sum is empty.
Therefore, $E_{m+1,n+m+1-\ell}=ny_{\ell-1}^{\prime}+e_{m+1,\ell-2}$ with $e_{m+1,\ell-2}\in\mathcal{R}\{Y_{\ell-2}\}$, and this conclude the proof. ∎

### v2

@@ 3.1 Formal differential operators
Then $\mathcal{R}$ is a differential ring whose ring of constants equals $\mathbf{C}$. We call $\mathcal{R}[\partial]$ the ring of formal differential operators in the differential variables $U$ and its elements formal differential operators.
Definition 3.6.
(Weight function) We define a weight function $w$ on $\mathcal{R}$ given by $w(u_{\ell})=\ell$, and
$w(u_{\ell}^{(k)})=\ell+{k},\ k=1,2,\dots.$ (6)
As in [25, Proposition 4.2], we call an operator $P\in\mathcal{R}[\partial]$ homogeneous of weight $r$ with respect to $w$ if the coefficient of $\partial^{i}$ in $P$ has weight rir-i. This is equivalent to extending $w$ to $\mathcal{R}[\partial]$ by setting $w(\partial)=1$. From the homogeneity of the commutation rule $\partial p=p\partial+\partial(p)$, for every $p\in\mathcal{R}$ it follows that the product of two operators that are homogeneous of weights $r$ and $s$ is homogeneous of weight $r+s$.
@@
which is homogeneous of weight $n$.
Remark 3.7.
An $n$-th root of $L$ in $\mathcal{R}((\partial^{-1}))$ is an order one pseudo-differential operator $\xi\partial+q_{0}+q_{-1}\partial^{-1}+\cdots$, where $\xi$ is an $n$-th root of unity. We work with the unique $Q$ starting with $\partial$ such that $Q^{n}=L$. The coefficient of the term of order n1n-1 of $Q^{n}$ is $nq_{0}$. Since $L$ is in normal form, then the term of order n1n-1 of $L$ is zero, implying that $q_{0}=0$. Thus $Q$ is also in normal form
$Q=\partial+q_{-1}\partial^{-1}+\cdots.$ (8)
The next theorem is our reformulation of [1, Proposition 2.4], which establishes the foundation of the contributions of this paper.
@@
We say that a pseudo-differential operator in $\mathcal{R}((\partial^{-1}))$ is homogeneous of weight $r$ if the coefficient of $\partial^{i}$ is homogeneous of weight rir-i, with $i\in\mathbb{Z}$. In $\mathcal{R}((\partial^{-1}))$, it also holds that the product of two operators that are homogeneous of weights $r$ and $s$ is homogeneous of weight $r+s$. It follows immediately that if a pseudo-differential operator is homogeneous of weight $r$ then its positive part is also homogeneous of weight $r$.
By [25, Proposition 4.2], the unique monic $n$-th root $Q=L^{1/n}$ of $L$ is a pseudo-differential operator in $\mathcal{R}((\partial^{-1}))$ and it is homogeneous of weight $1$. Therefore $Q^{j}$ is homogeneous of weight $j$ and we conclude that  $P_{j}=(Q^{j})_{+}$ is homogeneous of weight $j$. Since $L$ is monic and in normal form, then by Remark 3.7 so is $Q$, which implies that each $P_{j}$ is monic and in normal form.
To prove its uniqueness, let us consider any other $\mathbf{C}$-basis $\mathcal{T}=\{T_{j}\mid j\in\mathbb{N}\}$ of $W(L)$ satisfying the homogeneity condition. It is necessary that $T_{j}=c_{j}P_{j}$, with $c_{j}\in\mathbf{C}$, since otherwise it would not be homogeneous anymore. If we require every element of $\mathcal{T}$ to be monic, we conclude that $c_{j}=1$ for all $j\in\mathbb{N}$, proving $\mathcal{T}=\mathcal{H}(L)$. ∎
@@
is a $\mathbf{C}$-basis of $W_{M}(L)$. Our goal is to compute the basis $\mathcal{H}_{M}(L)$.
@@
Let us illustrate the homogeneous almost commuting basis in the case $n=3$, that leads to the Boussinesq hierarchy, which will be studied in Section 5.1. Consider a formal third-order differential operator
@@
Its monic cubic root is the pseudo-differential operator in $\mathcal{R}((\partial^{-1}))$
@@
The differential operator $P_{j}$ defined in (13) is a monic operator of order $j$ whose coefficient in $\partial^{j-1}$ is 0. For $j\equiv 0\,(\mod\ 3)$ observe that $P_{j}=L_{3}^{j}$.
In Section 4, we provide a deterministic algorithm to compute the sequence  $\{P_{j}\}$ defined in (9) up to a fixed bound $M$, for any formal differential operator of order $n$ as in (7). Our philosophy is to avoid computations in the ring of pseudo-differential operators $\mathcal{R}((\partial^{-1}))$, since the formal implementation of $\mathcal{R}((\partial^{-1}))$ in a computational algebra system is not available so far.
@@ 4 Basis of almost-commuting ODOs
Definition 4.1.
(Extended weight function) Let us define a new set of differential variables $Y=\{y_{2},\dots,y_{M}\}$ over $\mathcal{R}$. Let $Y_{m}$ denote $\{y_{2},\ldots,y_{m}\}$ for $m=2,\ldots,M$. Extending the derivation in $\mathcal{R}$ formally by
@@
which is homogeneous of weight $m$. Consequently, the Lie bracket of $L$ and $\tilde{P}_{m}$ is homogeneous of weight $n+m$ in $\mathbb{Q}\{Y_{m}\}[\partial]$. By Lemma 2.1, the Lie bracket $L,\tilde{P}_{m}]$ provides a formal differential operator of order less than or equal to n+m3n+m-3 since both operators are in normal form. The next equation will be of importance in the remaining parts of this paper:
$L,\tilde{P}_{m}]=E_{m,0}(U,Y_{m})+E_{m,1}(U,Y_{m})\partial+\ldots+E_{m,n+m-3}(U,Y_{m})\partial^{n+m-3},$ (16)
@@
$\epsilon_{Z}:\mathcal{R}\{Y_{m}\}\rightarrow\mathcal{R}\ ,\ {\epsilon_{Z}(y_{\ell}^{(k)})=q_{\ell}^{(k)}},$ (18)
and denote $f(U,Z):=\epsilon_{Z}(f(U,Y_{m}))$ . Abusing notation, we define the extension of the previous morphism to the corresponding rings of differential operators
@@
The next theorem explains that solving the system $\mathbb{E}_{m,n}$ in (17) for the set variables $Y_{m}$ is equivalent to finding the homogeneous basis $\mathcal{H}(L)$ of $W(L)$ as in Theorem 3.8. Moreover, by considering weights, the solution $Z=\{q_{2},\dots,q_{m}\}$ is a unique set of differential polynomials in $\mathcal{R}=\mathbf{C}\{U\}$.
@@
Consider a formal differential operator $L$ of order $n$ in normal form as in (7). Let us fix a positive integer $m\geq 2$, and the homogeneous operator $P_{m}$ in the homogeneous basis $\mathcal{H}(L)$ of $W(L)$ (see Theorem 3.8). Let us consider $Z=\{q_{2},\dots,q_{m}\}\subset\mathcal{R}$ with weights $w(q_{\ell})=\ell$. The following statements are equivalent:
@@
    $Z$ is a solution of system $\mathbb{E}_{m,n}$ in (17) w.r.t $Y_{m}$, that is
@@
The result follows from the next observation
$L,\epsilon_{Z}(\tilde{P}_{m})]=\epsilon_{Z}([L,\tilde{P}_{m}])=\sum_{j=3}^{n+m}E_{m,n+m-j}(U,Z)\partial^{n+m-j}.$ (20)
If we assume 1, since $L,P_{m}]$ is an operator of order n2n-2, then $Z$ is a solution of system (17). Now assume 2. Then, by (20), $\epsilon_{Z}(\tilde{P}_{m})$ almost commutes with $L$ and it is an homogeneous operator of weight $m$, by the weight condition on $q_{\ell}$. Therefore, by the uniqueness in Theorem 3.8, the operator $\epsilon_{Z}(\tilde{P}_{m})$ equals $P_{m}$. ∎
@@
For each $n,m\in\mathbb{N}$, $n,m\geq 2$, the system $\mathbb{E}_{m,n}$ has a unique solution $Z=\{q_{2},\dots,q_{m}\}\in\mathbb{Q}\{U\}$ in the variables $Y_{m}$, with weights $w(q_{\ell})=\ell$. We call $Z$ the weighted solution of $\mathbb{E}_{m,n}$
@@ 2 LB $\leftarrow[L,\tilde{P}_{m}]$;
3 $\mathbb{E}_{m,n}$ $\leftarrow\left\{\operatorname{coeff}(\texttt{LB},\partial^{i})\ :\ i\in\{n-1,\ldots,n+m-3\}\right\}$;
@@ 4 Compute the unique weighted solution $Z$ of system $\mathbb{E}_{m,n}$ in (17);
The previous results induce Algorithm  1 to compute the almost commuting operator $P_{m}$ of the basis $\mathcal{H}(L)$. Iterating Algorithm 1 for $m=2,\ldots,M$ and setting $P_{1}=\partial$, we can compute the homogeneous basis of $W_{M}(L)$. We will discuss step 4 of this algorithm in the remaining parts of this section.
@@ 4.1 Triangular differential system
is a triangular system in the set of differential variables $Y_{m}=\{y_{2},\ldots,y_{m}\}$. In addition, each differential polynomial $E_{m,m+n-i}(U,Y_{m})$ will be proved to be homogeneous of weight $i$.
@@
[…]
This induces the following recursive formula to compute the Lie-brackets $L,\tilde{P}_{m}]$.
@@
Let $L$ be as in (7). Then, for $\ell\geq 1$
$L,\tilde{P}_{\ell+1}]=[L,\tilde{P}_{\ell}]\partial+\tilde{P}_{\ell}[L,\partial]+[L,y_{\ell+1}].$ (22)
@@
We can see that the first summand adds the coefficients of  $\partial^{i}$ in $L,\tilde{P}_{\ell}$] as coefficients of $\partial^{i+1}$ in $L,\tilde{P}_{\ell+1}]$. The second summand can be easily computed as a product of differential operators (see Lemma 4.4). Finally, the third summand has also a very generic formula (see Lemma 4.5) and is the only summand that involves the differential variable $y_{\ell+1}$.
@@
    $L,\partial]=-\left(u_{n}^{\prime}+u_{n-1}^{\prime}\partial+\ldots+u_{2}^{\prime}\partial^{n-2}\right)$. Therefore, it is a homogeneous operator of weight $n+1$.
@@
    For $\ell\geq 2$, the differential operator $Q_{\ell}:=\tilde{P}_{\ell}[L,\partial]$ has order $n+\ell-2$, and it is a homogeneous operator of weight $\ell+n+1$. In particular, the coefficients of $Q_{\ell}$ only contain the differential variables $y_{2},\ldots,y_{\ell}$ with no derivatives. Moreover, the coefficient of $y_{\ell-i}\partial^{k}$ in $Q_{\ell}$ is 0, for k>n+i2k>n+i-2.
@@
The formula in 1 follows from the commutation rule $\partial u_{k}=u_{k}\partial+u_{k}^{\prime}$. Next, for proving 2, we first observe that $\operatorname{ord}(Q_{\ell})=\ell+n-2$, and, since $\tilde{P}_{\ell}$ is homogeneous of weight $\ell$ and $L,\partial]$ is homogeneous of weight $n+1$, then $Q_{\ell}$ is also homogeneous of weight $\ell+n+1$. Moreover, we can compute
@@
$=-u_{2}^{\prime}\partial^{n+\ell-2}-\left(\ell u_{2}^{\prime\prime}+u_{3}^{\prime}\right)\partial^{n+\ell-3}-\sum_{i=0}^{\ell-2}y_{\ell-i}H_{n+i-2}+\left(\textrm{l.o.t.~in }\mathbf{C}\{U\}[\partial]\right),$
with $H_{n+i-2}$ a ordinary differential operator in $\mathbf{C}\{U\}[\partial]$ of order at most n+i2n+i-2. Then, the result follows. ∎
@@
Let $\ell$ be a positive integer. Then $L,y_{\ell+1}]$ is a homogeneous differential operator of order n1n-1, weight $n+\ell+1$, and leading coefficient $ny_{\ell+1}^{\prime}$.
@@
Computing the commutator $L,y_{\ell+1}]=Ly_{\ell+1}-y_{\ell+1}L$, we obtain
$[L,y_{\ell+1}]$ $=\displaystyle\left(\partial^{n}+\sum_{i=0}^{n-2}u_{n-i}\partial^{i}\right)y_{\ell+1}-\left(y_{\ell+1}\partial^{n}+\sum_{i=0}^{n-2}y_{\ell+1}u_{n-i}\partial^{i}\right)$
$=\sum_{k=0}^{n-1}\binom{n}{k}y_{\ell+1}^{(n-k)}\partial^{k}+\sum_{i=0}^{n-2}\sum_{k=0}^{i}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\partial^{k}-\sum_{k=0}^{n-2}y_{\ell+1}u_{n-k}\partial^{k}$
$=\sum_{k=0}^{n-1}\binom{n}{k}y_{\ell+1}^{(n-k)}\partial^{k}+\sum_{k=0}^{n-2}\sum_{i=k}^{n-2}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\partial^{k}-\sum_{k=0}^{n-2}y_{\ell+1}u_{n-k}\partial^{k}$
$=ny_{\ell+1}^{\prime}\partial^{n-1}+\displaystyle\sum_{k=0}^{n-2}\left[\binom{n}{k}y_{\ell+1}^{(n-k)}+\sum_{i=k+1}^{n-2}\binom{i}{k}u_{n-i}y_{\ell+1}^{(i-k)}\right]\partial^{k}$
$=ny_{\ell+1}^{\prime}\partial^{n-1}+\textrm{l.o.t.}$
@@
We prove next the triangularity of the system $\mathbb{E}_{m,n}$ defined in (17). We provide algorithms to solve the system $\mathbb{E}_{m,n}$ in Section 4.2.
@@
where $e_{m,i-2}\in\mathcal{R}\{Y_{i-2}\}$ is homogeneous of weight $i$, for $i\geq 3$ (setting $Y_{1}=\emptyset$). More precisely, $E_{m,m+n-3}=ny^{\prime}_{2}-mu^{\prime}_{2}$.
@@
We proceed by induction on $m$. First, for $m=2$ we use Lemmas 4.3 and 4.4 to we obtain
$L,\tilde{P}_{2}]=-2\sum_{p=1}^{n-1}u^{\prime}_{p+1}\partial^{n-p}-\sum_{p=2}^{n}u^{\prime\prime}_{p}\partial^{n-p}+[L,y_{2}].$
@@
Now, assume the theorem holds for $m\geq 2$ and let us prove the case $m+1$. Using the recursive structure of $L,\tilde{P}_{m+1}]$ from Lemma 4.3, formula (15) and Lemma 4.4, the following recursion holds
$L,\tilde{P}_{m+1}]=\!\sum_{j=0}^{n+m-3}\!\!E_{m,j}\partial^{j+1}-\Bigg{(}\partial^{m}+\sum_{j=0}^{m-2}y_{m-j}\partial^{j}\Bigg{)}\Bigg{(}\sum_{p=2}^{n}u^{\prime}_{p}\partial^{n-p}\Bigg{)}+[L,y_{m+1}].$
We use the following notations to simplify the previous identity:
$\left\{\begin{aligned} A_{j}&:=\partial^{j}\displaystyle\sum_{p=2}^{n}u^{\prime}_{p}\partial^{n-p}\in\mathcal{R}[\partial]\ ,\textrm{ of order }n+j-2,\\ \alpha_{j,\ell}&:=\operatorname{coeff}(A_{j},\partial^{\ell})\in\mathcal{R},\vskip 10.00002pt\\ R_{m+1}&:=[L,y_{m+1}]-ny^{\prime}_{m+1}\partial^{n-1}\in\mathcal{R}\{y_{m+1}\}[\partial]\text{ of order }n-2.\end{aligned}\right$.
@@
$L,\tilde{P}_{m+1}]=\sum_{j=0}^{n+m-3}E_{m,j}\partial^{j+1}-A_{m}-\sum_{j=0}^{m-2}y_{m-j}A_{j}+ny^{\prime}_{m+1}\partial^{n-1}+R_{m+1}.$
@@
Since $A_{j}$ has order n+j2n+j-2, it is clear that $\alpha_{j,\ell}=0$ for $\ell>n+j-2$, so, in particular, $\alpha_{j,n+m-2}=0$ for $j<m$. Moreover, we can observe that $\alpha_{j,n+j-2}=u_{2}^{\prime}$. Hence, using the recursion for $E_{m,n+m-3}$ we get:
@@
Finally, for all the intermediate equations, $4\leq\ell\leq m+1$, we observe that
@@
But, by the induction hypothesis, the first summand equals $E_{m,n+m-\ell}=ny_{\ell-1}^{\prime}+e_{m,\ell-2}$, where $e_{m,\ell-2}$ is in $\mathcal{R}\{Y_{i-2}\}$. The second summand is in $\mathcal{R}$, and we can observe that the third summand involve the variables $y_{k}$ for $k=2,\ldots,\ell-3$. Observe that in the case of $\ell=4$ this sum is empty.
Therefore, $E_{m+1,n+m+1-\ell}=ny_{\ell-1}^{\prime}+e_{m+1,\ell-2}$ with $e_{m+1,\ell-2}\in\mathcal{R}\{Y_{\ell-2}\}$, and this concludes the proof. ∎